Manny is pushing a 5.0 kg box across a level floor with a force of 50.0N. The force he applies is downward and to the right at 250. The coefficient of kinetic friction between the box and the floor is 0.22. Find the acceleration of the box.
1
Expert's answer
2020-11-16T07:38:42-0500

Answer

Force is applied downward 25° to the right.

So

"F_{net}=ma"

And

"F_{net}=Fcos\theta-\mu (mg+Fsin\theta)"

By putting the value


"F_{net}=50\times cos25°-0.22\times(5\times9.8+50sin25°)"

"F=30N"

And

"30=5a\\
So\\
a=\frac{30}{5}=6m/s^2"


So acceleration of box is 6m/s^2.